# Turning English-centric LLMs Into Polyglots: How Much Multilinguality Is Needed?

### v1

@@ E.4 XNLI
Figure 14 shows the results of Falcon 7b on XNLI. Similar to the results attained with Llama 2 7b (see Figure 5), we observe no significant differences in performance given different degrees of multilingual instruction tuning.
E.5 Discussion
Overall, these experiments indicate that Falcon’s multilingual capabilities are considerably narrower than Llama 2’s. We hypothesise that this may be a result of more stringent filtering of web-scraped pretraining data Penedo et al. (2023b), possibly resulting in less accidental contamination. Interestingly though, according to our estimated statistics based on the sample of the RefinedWeb corpus (Table 1), which makes up a large portion of Falcon’s pretraining data, Russian appears as frequently as Chinese and Swedish. However, we find that Russian texts are, on average, three times longer when tokenized with Falcon’s tokenizer compared to Llama 2’s. This indicates that, despite having a much larger vocabulary (65k for Falcon vs. 32k for Llama 2), Cyrillic-scripted languages such as Russian and Bulgarian are outside of the model’s intended domain and performance on these languages is inherently limited.
@@ E.5 Discussion
Figure 11: Average helpfulness of chat responses from Falcon 7b with incremental multilingual instruction tuning. Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
@@
Figure 12: X-QuAD results for Falcon 7b with incremental multilingual instruction tuning. Results are shown for both multilingual prompting (en:xx) and monolingual prompting (xx:xx). Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
@@
Figure 13: X-CSQA results for Falcon 7b with incremental multilingual instruction tuning. Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
@@
Figure 14: XNLI results for Falcon 7b with incremental multilingual instruction tuning. Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
@@
You are an expert language evaluator. You are evaluating a response that has been submitted for a particular task, using a specific set of standards. Below is the data: [BEGIN DATA] *** [Task]: {{task}} *** [Submission]: {{submission}} *** [Criterion]: helpfulness: ‘‘1’’: ‘‘Not helpful - The generated text is completely irrelevant, unclear, or incomplete. It does not provide any useful information to the user.’’ ‘‘2’’: ‘‘Somewhat helpful - The generated text has some relevance to the user’s question, but it may be unclear or incomplete. It provides only partial information, or the information provided may not be useful for the user’s needs.’’ ‘‘3’’: ‘‘Moderately helpful - The generated text is relevant to the user’s question, and it provides a clear and complete answer. However, it may lack detail or explanation that would be helpful for the user.’’ ‘‘4’’: ‘‘Helpful - The generated text is quite relevant to the user’s question, and it provides a clear, complete, and detailed answer. It offers additional information or explanations that are useful for the user. However, some of the points of the response are somewhat repetitive or could be combined for greater clarity and concision.’’ ‘‘5’’: ‘‘Very helpful - The generated text is highly relevant to the user’s question, and it provides a clear, complete, and detailed answer. It offers additional information, explanations, or analogies that are not only useful but also insightful and valuable to the user. However, the structured of the response is not well-organized and there is no clear progression or logical sequence of different points in the response.’’ ‘‘6’’: ‘‘Highly helpful - The generated text provides a clear, complete, and detailed answer. It offers additional information or explanations that are not only useful but also insightful and valuable to the user. The response is also in a logical and easy-to-follow manner by explicitly using headings, bullet points, or numbered lists to break up the information and make it easier to read.’’ *** [END DATA] Does the submission meet the criterion? First, write out in a step by step manner your reasoning about the criterion to be sure that your conclusion is correct. Avoid simply stating the correct answers at the outset. Provide your response as a RFC8259 compliant JSON following this schema: {‘‘reasoning’’: str, ‘‘score’’: int}
@@
Figure 16: Prompt template used to query the LLM judge in order to assess the helpfulness of an open-ended chat response. This prompt closely follows the one used by Zhou et al. (2023) but extends it by specifying the return value as a valid JSON-like object.
Prompt:
@@
### Human: Answer the question from the given passage. Your answer should be directly extracted from the passage, and it should be a single entity, name, or number, not a sentence.
@@
Reference:
zwei
System:
two".
@@
Table 2: Example input prompt, reference and model response for the XQuAD task. Monospaced font represents the prompt template used for each test instance. Here, we show the multilingual prompting strategy (en:xx), where the task instruction and response prefix is provided English, while the context passage and question are in the target language (in this case, German). A correct answer in this setting requires the model to continue the response by switching into the target language, which it fails to do in this particular example.
Prompt:
### Human: Dir wird eine Frage vorgelegt, die mehrere Antwortmöglichkeiten hat. Wähle die am besten geeignete Option aus ‘‘A’’, ‘‘B’’, ‘‘C’’, ‘‘D’’ oder ‘‘E’’ basierend auf deinem gesunden Menschenverstand. Bitte gebe deine Antwort in Form eines einzelnen Buchstabens in Anführungszeichen an.
@@
Reference:
E
System:
E”
Table 3: Example of the X-CSQA task. Monospaced font represents the prompt template used for each test instance. Here, the prompt strategy is monolingual (xx:xx) with the task instruction provided in the same language as the test instance (in this case, German). A correct answer requires the model to output the letter corresponding to the ground-truth answer.

### v2

@@ F.5 XNLI
Figure 16 shows the results of Falcon 7b on XNLI. Similar to the results attained with Llama 2 7b (cf. Figure 6), we observe no significant differences in performance given different degrees of multilingual instruction tuning.
F.6 Discussion
These additional experiments using Falcon 7b provide further support for our main findings: instruction tuning with as few as two languages significantly improves cross-lingual generalisation, enhancing performance in open-ended single-turn dialogue for some non-English languages. However, strikingly, multilingual capabilities of Falcon are considerably narrower than that of Llama 2. This is also suggested by the statistics provided in Table 1, which show a higher concentration of English and much lower proportions of non-English languages compared with both Llama 2 and GPT-3. We suspect that this may be a result of stringent filtering of web-scraped pretraining data performed in producing the RefinedWeb corpus used to train the Falcon models Penedo et al. (2023), which not only reduces the risk of potentially accidental contamination Blevins and Zettlemoyer (2022) but also language coverage. That said, the RefinedWeb corpus comprises only part of Falcon’s actual training corpus. Almazrouei et al. (2023) note that the final corpus also contains additional data drawn from curated sources including a European-focused (multilingual) Common-Crawl dataset, which could explain the strong performance on some major European languages. In light of this, we suspect that additional Chinese data is also included in Falcon’s training corpus, allowing for relatively strong performance in Chinese on the single-turn dialogue task.
@@ F.6 Discussion
Figure 12: Average helpfulness of single-turn dialogue responses from Falcon 7b given incremental multilingual instruction tuning. Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
@@
Figure 13: SARI weighted by IO language agreement for sentence simplification with Falcon 7b given incremental multilingual instruction tuning. Results are shown for both cross-lingual prompting (en:xx) and monolingual prompting (xx:xx) Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
@@
Figure 14: XQuAD results for Falcon 7b given incremental multilingual instruction tuning. Results are shown for both cross-lingual prompting (en:xx) and monolingual prompting (xx:xx). Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
@@
Figure 15: X-CSQA results for Falcon 7b given incremental multilingual instruction tuning. Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
Figure 16: XNLI results for Falcon 7b given incremental multilingual instruction tuning. Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
Appendix G Results with Llama 3 8b
In §5.2 we observed that scaling the base model from 7b to 70b parameters reduces the gap in performance between English and non-English target languages with regard to the single-turn dialogue task. In this section, we evaluate Llama 3 8b MetaAI (2024), which builds on Llama 2 by scaling up the pretraining data from 2 trillion tokens to 15 trillion tokens. In Addition to a much larger pretraining corpus, Llama 3 is trained with a vocabulary of 128k tokens (vs. 32k for Llama 2), bringing it much closer to traditional multilingual models which typically contain roughly 250k tokens Le Scao et al. (2023); Xue et al. (2021); Conneau et al. (2020). While the exact language distribution of the model’s pretraining data is not publicly known, it is reasonable to expect that these enhancements could lead to improvements in the model’s non-English capabilities.
Figure 17: Single-turn dialogue performance of Llama 3 8b given incremental multilingual instruction tuning. Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
Figure 17 shows the performance of Llama 3 8b, which closely resembles that of Llama 2 70b across most non-English languages, even outperforming it in languages like Vietnamese, Greek, and Hindi. Most notably, we observe that our main findings still hold: multilingual instruction tuning with as few as two languages is both necessary and sufficient to elicit cross-lingual generalisation. Again, performance typically plateaus thereafter. Given the improved performance on languages like Greek and Hindi that are low-resource under Llama 2’s pretraining data distribution, we suspect that their representation in the pretraining data for Llama 3 is substantially higher. Still, performance on Icelandic suggests that this language may remain in the low-resource category under this model.
Appendix H Ablation Experiments
In Section 3.2, we constructed Multi-$i$ instruction tuning datasets of the same size by replacing a fixed number of English examples with examples from a new language. As a result, the proportion of non-English examples used for finetuning increases along with the number of languages. Table 3 shows the exact makeup of each these datasets and how they relate to one another. To rule out potential confounding factors between the proportion of non-English examples and the diversity of languages, we conduct ablation experiments in which we keep one variable fixed while incrementing the other.
Dataset Languages % NE Total
Mono en 0.00% 3,200
Multi-2 en, es 6.25% 3,200
Multi-3 en, es, ru 12.50% 3,200
Multi-4 en, es, ru, de 18.75% 3,200
Multi-5 en, es, ru, de, zh 25.00% 3,200
Multi-6 en, es, ru, de, zh, fr 31.25% 3,200
Table 3: Makeup of the incremental multilingual instruction-tuning datasets.
[…]
(b) Impact of incrementing the total number of non-English examples while keeping the number of languages fixed at three (English, Spanish, and Russian). Here, Russian is seen during finetuning, while German and Chinese are in the 0-shot setting.
Figure 18: Average helpfulness of single-turn dialogue responses from Llama 2 7b given multilingual instruction tuning on a fixed budget (limiting either the number of non-English instances 18(a) or the number of languages 18(b)). Error bars show a confidence interval of 95%.
Figure 18(a) shows that when using a fixed budget of 400 non-English examples (which equates to 12.5% of the finetuning data) and incrementing only the number of languages within this budget, performance increases consistently as each new language is added, closely reflecting the results of our main experiments. In contrast, Figure 18(b) shows that when training with a fixed number languages (e.g., Multi-2) and incrementing the number of non-English examples (in this case, Spanish), performance tends to remain lower and generally shows less improvement. These results underscore our main findings discussed in §4.1 and §6.
@@ Appendix H Ablation Experiments
You are an expert language evaluator.
You are evaluating a response that has been submitted for a particular task, using a specific set of standards. Below is the data:
[BEGIN DATA]
[Task]: {{task}}
[Submission]: {{submission}}
[Criterion]: helpfulness:
‘‘1’’: ‘‘Not helpful - The generated text is completely irrelevant, unclear, or incomplete. It does not provide any useful information to the user.’’
‘‘2’’: ‘‘Somewhat helpful - The generated text has some relevance to the user’s question, but it may be unclear or incomplete. It provides only partial information, or the information provided may not be useful for the user’s needs.’’
‘‘3’’: ‘‘Moderately helpful - The generated text is relevant to the user’s question, and it provides a clear and complete answer. However, it may lack detail or explanation that would be helpful for the user.’’
‘‘4’’: ‘‘Helpful - The generated text is quite relevant to the user’s question, and it provides a clear, complete, and detailed answer. It offers additional information or explanations that are useful for the user. However, some of the points of the response are somewhat repetitive or could be combined for greater clarity and concision.’’
‘‘5’’: ‘‘Very helpful - The generated text is highly relevant to the user’s question, and it provides a clear, complete, and detailed answer. It offers additional information, explanations, or analogies that are not only useful but also insightful and valuable to the user. However, the structured of the response is not well-organized and there is no clear progression or logical sequence of different points in the response.’’
‘‘6’’: ‘‘Highly helpful - The generated text provides a clear, complete, and detailed answer. It offers additional information or explanations that are not only useful but also insightful and valuable to the user. The response is also in a logical and easy-to-follow manner by explicitly using headings, bullet points, or numbered lists to break up the information and make it easier to read.’’
[END DATA]
Does the submission meet the criterion? First, write out in a step by step manner your reasoning about the criterion to be sure that your conclusion is correct. Avoid simply stating the correct answers at the outset. Provide your response as a RFC8259 compliant JSON following this schema:
{‘‘reasoning’’: str, ‘‘score’’: int}
@@
Figure 20: Prompt template used to query the LLM judge in order to assess the helpfulness of a single-turn dialogue response. This prompt closely follows the one used by Zhou et al. (2023) but extends it by specifying the return value as a valid JSON-like object.
Prompt: ### Human: Please rewrite the following complex sentence in order to make it easier to understand. You can do so by replacing complex words with simpler synonyms (i.e. paraphrasing), deleting unimportant information (i.e. compression), and/or splitting a long complex sentence into several simpler ones. The final simplified sentence needs to be grammatical, fluent, and retain the main ideas of its original counterpart without altering its meaning.
Complex sentence: Sie gelten als eiskalte Fress-Maschinen.
### Assistant:
Reference: Sie gelten als Fress-Maschinen.
System: They are seen as cold-blooded eaters.
Table 4: An example of the 0-shot sentence simplification task. Monospaced font represents the prompt template used, adapted from instructions provided to crowdworkers for the creation of the ASSET dataset Alva-Manchego et al. (2020). This example shows the cross-lingual prompt strategy (en:xx), where the task instruction is provided in English but the source sentence is in German. Ideally, the system language should match the source sentence, which it fails to do in this case.
@@
Prompt: ### Human: Answer the question from the given passage. Your answer should be directly extracted from the passage, and it should be a single entity, name, or number, not a sentence.
@@
Reference: zwei
System: two".
@@
Table 5: Example input prompt, reference, and model response for the 0-shot extractive QA task using XQuAD. Monospaced font represents the prompt template used for each test instance. Here, we show the cross-lingual prompting strategy (en:xx), where the task instruction and response prefix is provided English, while the context passage and question are in the target language (in this case, German). A correct answer in this setting requires the model to continue the response by switching into the target language, which it fails to do in this particular example.
Prompt: ### Human: Dir wird eine Frage vorgelegt, die mehrere Antwortmöglichkeiten hat. Wähle die am besten geeignete Option aus ‘‘A’’, ‘‘B’’, ‘‘C’’, ‘‘D’’ oder ‘‘E’’ basierend auf deinem gesunden Menschenverstand. Bitte gebe deine Antwort in Form eines einzelnen Buchstabens in Anführungszeichen an.
@@
Reference: E
System: E”
Table 6: Example of the commonsense reasoning task using X-CSQA. Monospaced font represents the prompt template used for each test instance. Here, the prompt strategy is monolingual (xx:xx) with the task instruction provided in the same language as the test instance (in this case, German). A correct answer requires the model to output the letter corresponding to the ground-truth answer.
XNLI prompt template: {Sentence 1}, right? {label}, {Sentence 2}
Label Prompt query NLL
Entailment In measuring effectiveness, perfection is unattainable., right? Yes, You can never be perfect. -113.94
Neutral In measuring effectiveness, perfection is unattainable., right? Also, You can never be perfect. -116.79
Contradiction In measuring effectiveness, perfection is unattainable., right? No, You can never be perfect. -114.01
Table 7: Example prompt queries for XNLI given the predefined template used in the LM Evaluation Harness Gao et al. (2023) and the three possible labels. Underlined words are the language-specific connectors corresponding to the three possible labels. The final answer is the sequence with the highest negative log-likelihood (NLL) according to the model. In this case, Llama 2 7b Mono correctly identifies the relationship between the two sentences as entailment.